# Score-Based Metropolis-Hastings Algorithms

### v1

@@ Abstract
In this paper, we introduce a new approach for integrating score-based models with the Metropolis-Hastings algorithm. While traditional score-based diffusion models excel in accurately learning the score function from data points, they lack an energy function, making the Metropolis-Hastings adjustment step inaccessible. Consequently, the unadjusted Langevin algorithm is often used for sampling using estimated score functions. The lack of an energy function then prevents the application of the Metropolis-adjusted Langevin algorithm and other Metropolis-Hastings methods, limiting the wealth of other algorithms developed that use acceptance functions. We address this limitation by introducing a new loss function based on the detailed balance condition, allowing the estimation of the Metropolis-Hastings acceptance probabilities given a learned score function. We demonstrate the effectiveness of the proposed method for various scenarios, including sampling from heavy-tail distributions.
@@ 1 Introduction
Sampling from probability distributions with only access to samples is a longstanding problem with significant attention devoted to it in the statistics and machine learning communities (Goodfellow et al., 2020; Kingma, 2013; Ho et al., 2020; Dinh et al., 2016; Song et al., 2021). A wealth of literature and research is devoted to studying efficient methods for sampling from densities where only a function proportional to the density is known (Hastings, 1970; Wang et al., 2024). A particular instance of this is the Metropolis-Hastings (MH) method, which provides a condition for accepting or rejecting new samples based on knowledge of the function proportional to the density (Robert et al., 2004). The MH algorithm builds a Markov chain based on a proposal distribution and an acceptance function and its convergence is guaranteed when the process satisfies the detailed balance condition. The success of this framework led to a significant amount of research effort devoted to algorithms that make use of the detailed balance criterion. Various MH algorithms have been introduced, defined by their choice of proposal distributions, such as the Random Walk Metropolis (RW) algorithm (Metropolis et al., 1953), the Metropolis Adjusted Langevin Algorithm (MALA) (Roberts and Tweedie, 1996; Roberts and Stramer, 2002), and the preconditioned Crank-Nicolson (pCN) algorithm (Hairer et al., 2014).
While these methods are theoretically well-grounded and provide strong empirical performance, they require knowledge of the unnormalized density, which may be difficult to obtain when only given samples. In recent years, score-based models have gained prominence due to their ability to learn the score function, i.e., the gradient of the log-density, directly from data samples (Song and Ermon, 2019; Song et al., 2020). Consequently, score-based sampling methods, such as the unadjusted Langevin algorithm, have primarily been employed to leverage the learned score function to sample from the target distribution (Grenander and Miller, 1994; Roberts and Tweedie, 1996).
@@
With only knowledge of the score function, the research devoted to sampling using MH criteria is not used within the context of score-based generative modeling. This work proposes a unifying framework that bridges this gap, enabling the application of MH algorithms using only samples and a learned score function. Our analysis is motivated by two questions:
@@
1. Can the acceptance function of an MH algorithm be estimated solely based on the knowledge of an estimated score function? 2. Does including an MH adjustment step improve sample quality for score-based models?
@@
We answer these questions affirmatively by developing an approach to estimate an acceptance probability function from data and evaluating its effectiveness by comparing to ULA algorithms.
@@ 2 Related Work
A significant area of related work begins with estimating a score function for the data and then finding efficient ways of sampling using this estimated score function. While majority of such methods focus on designing sampling schemes that converge to the target distribution faster (e.g. through different numerical integrators), we consider our work as a parallel approach which uses an acceptance function to accept or reject different samples. This work lies at the intersection of the literature on score-based models and the Markov Chain Monte Carlo (MCMC) techniques that use Metropolis-Hastings acceptance functions. We then review recent works on score-based models and Metropolis-Hastings algorithms.
@@
As mentioned within the introduction, many of the empirically successful sampling techniques involve sampling using an estimated score function (Song and Ermon, 2019; Song et al., 2021). These methods are trained using the score loss function by minimizing the Fisher divergence between the true and the predicted score (Hyvärinen, 2005; Vincent, 2011). A significant research effort is concerned with understanding the efficacy of these sampling algorithms, with some papers investigating the asymptotic error of score-based diffusion models, as in (Chen et al., 2023; Zhang et al., 2024). Different sampling schemes have been proposed to improve sampling speed and quality (Dockhorn et al., 2021; Lu et al., 2022; Chen et al., 2024). These methods have relied heavily on various score-based samplers such as ULA or the reverse-time SDE of a diffusion process (Song et al., 2021). Notably, Sjöberg et al. (2023) describe a line integration method to compute the acceptance probabilities.
Metropolis-Hastings Methods.
Metropolis-Hastings (MH) algorithms are a class of MCMC algorithms that use an acceptance function to sample from target distributions (Metropolis et al., 1953; Hastings, 1970). These algorithms generate a sequence of samples by proposing a candidate state from a proposal distribution and accepting or rejecting the candidate based on an acceptance criterion that ensures convergence to the target distribution. The efficiency of MH algorithms heavily depends on the choice of the proposal distribution, and significant research effort has focused on designing efficient proposal mechanisms (Rosenthal et al., 2011; Song et al., 2017; Titsias and Dellaportas, 2019; Davies et al., 2023; Lew et al., 2023).
Recent developments in MH algorithms have explored the use of adaptive and state-dependent proposals to improve convergence rates and reduce autocorrelation in the generated samples (Andrieu and Thoms, 2008; Haario et al., 2001). These adaptive MH methods adjust the proposal distribution during the sampling process to better capture the geometry of the target distribution, leading to faster and more accurate sampling (Roberts and Rosenthal, 2009; Brooks et al., 2011; Hirt et al., 2021; Biron-Lattes et al., 2024).
(a) Original, weights $(0.8,0.2)$.
@@ 3 Score-Based Metropolis-Hastings
Let $X\sim p$ be a random variable, with support $\mathrm{supp}(X)=\mathcal{X}\subset\mathbb{R}^{d}$ and $p$ be its probability density function. We assume that $p$ is unknown, and only samples of $X$ are observed. Denote these observed samples by $\{x^{\left(i\right)}\}_{i=1}^{N}$, where $N$ is the number of observed samples.
The MH algorithm involves an acceptance function $a:\mathcal{X}\times\mathcal{X}\rightarrow[0,1]$, which defines the probability $a(x^{\prime},x)$ of transitioning from the current $x\in\mathcal{X}$ to a proposal $x^{\prime}\in\mathcal{X}$. The common choice of the acceptance function is given by:
@@
Note that for computing $a(x^{\prime},x_{t})$ in (1), a function proportional to $p(x)$ is needed. On the other hand, when sampling from $p(x)$, a common strategy is to estimate the score function $\nabla_{x}\log p(x)$ from data. In this work, we will assume that $\nabla_{x}\log p(x)$ is known through prior knowledge or estimated from the observed samples.
@@
Although the acceptance function is often chosen according to (1), under mild conditions, the sufficient condition for convergence to the target distribution is that the acceptance ratio satisfies the detailed balance condition. The detailed balance condition, also known as the Markov reversibility condition, is defined as:
@@
Acceptance functions satisfying the detailed balance condition allow the Markov chain to have the target distribution as its unique stationary distribution under some mild assumptions. In particular, (1) is a particular solution to this condition. Our goal is then to use the knowledge of the stationary condition in (2) such that our sampling scheme is informed by this regime.
Balance Matching.
To do this, we develop a loss function to estimate a valid acceptance function from only the score function based on the detailed balance condition given in (2). In the following proposition, we present the main observation motivating our method. We denote by $\nabla$ the gradient with respect to $x,x^{\prime}\in\mathbb{R}^{d}$, i.e., $\nabla=\nabla_{x,x^{\prime}}$. Thus, for a function $f$ with values in $\mathbb{R}$, the gradient $\nabla f$ has values in $\mathbb{R}^{2d}$.
Proposition 1.
If the gradients of the log acceptance ratio satisfy the following condition: for every $x,x^{\prime}\in\mathcal{X}$,
$\nabla\log\left(\frac{a(x^{\prime},x)}{a(x,x^{\prime})}\right)=\nabla\log\left(\frac{p(x^{\prime})}{p(x)}\right)+\nabla\log\left(\frac{q(x\mid x^{\prime})}{q(x^{\prime}\mid x)}\right)$ (3)
then the acceptance function $a(x^{\prime},x)$ must satisfy the equality:
$\frac{a(x,x^{\prime})}{a(x^{\prime},x)}=\frac{p(x^{\prime})q(x\mid x^{\prime})}{p(x)q(x^{\prime}\mid x)}.$
Proof.
Proof is provided in appendix C. ∎
It follows from Proposition 1 that matching the gradients of the left-hand side and right-hand side in (2) is a sufficient condition for verifying the detailed balance condition. In other words, if:
$\nabla\log a(x^{\prime},x)-\nabla\log a(x,x^{\prime})$
$=\nabla\log p(x^{\prime})-\nabla\log p(x)+\nabla\log q(x\mid x^{\prime})$
$\hskip 115.0pt-\nabla\log q(x^{\prime}\mid x),$
then the detailed balance condition is satisfied.
The acceptance function $a(x^{\prime},x)$ can be modeled using an expressive class of functions, such as neural networks, and this function can be optimized to satisfy the condition in (3). In this setup, the gradients $\nabla\log a(x^{\prime},x)$ and $\nabla\log a(x,x^{\prime})$ correspond to the model’s gradients, which can be efficiently computed via backpropagation when the model is differentiable. The terms $\nabla\log p(x)$ and $\nabla\log p(x^{\prime})$ represent the score functions at $x$ and $x^{\prime}$, which can be approximated with a score model. We can express these score functions as:
and, similarly,
MH with an Estimated Acceptance Functions.
In cases where the only a score function is available, we propose to estimate the acceptance function with balance matching using the estimated score function so that MH algorithms can be applied. We term this approach Score-Based Metropolis-Hastings. For the scope of this paper, we will instantiate our method on three representative MH algorithms, specified by their proposal functions: the Random Walk Metropolis-Hastings (RW), the Metropolis-Adjusted Langevin Algorithm (MALA), and the Preconditioned Crank-Nicolson (pCN) algorithm. We next briefly review these methods.
(a) Standard Random Walk Sampling.
(b) Score-based RW with true scores.
(c) Score-based RW with learned score.
Random Walk Metropolis-Hastings: In the RW algorithm, the proposal function is typically a Gaussian centered at the current state, i.e., $q(x^{\prime}\mid x)=\mathcal{N}\left(x,\sigma^{2}I\right)$, where $\sigma^{2}$ controls the step size of the random walk. This proposal is simple and symmetric, making it easy to implement.
Metropolis Adjusted Langevin Algorithm: MALA uses both the current state and the gradient of the log-probability. The proposal distribution is defined as $q(x^{\prime}\mid x)=\mathcal{N}\left(x+\frac{\varepsilon^{2}}{2}\nabla\log p(x),\varepsilon^{2}I\right)$, where $\varepsilon$ is a step size parameter. This approach leverages the local gradient information to propose moves that are more likely to be accepted.
Preconditioned Crank-Nicolson: The pCN algorithm modifies the standard Crank-Nicolson approach by introducing a preconditioning step. The proposal function for pCN is given by $x^{\prime}=\sqrt{1-\beta^{2}}x+\beta\xi$, where $\xi\sim\mathcal{N}(0,I)$ is a standard Gaussian noise term and $\beta$ controls the step size.
When the acceptance function is learned using balance matching, as defined in (3), we refer to these methods as Score RW, Score MALA, and Score pCN, respectively. While we illustrate our results on these three MH algorithms, we note that the proposed framework is general and can be applied to other proposal functions.
Slow Mixing.
There are two main motivations for our approach. The first is the challenge of applying traditional MH algorithms to generative modeling without requiring knowledge of the unnormalized density, while assuming that the score function can be accurately learned. The second motivation arises from limitations of the unadjusted Langevin algorithm (ULA), which, although commonly used for sampling from score-based models, does not guarantee convergence to the true distribution due to the impact of discretization.
Consequently, by incorporating an MH adjustment step, we address a key limitation observed in the ULA when sampling from distributions with multiple modes, particularly when these modes are separated by regions of low density. Specifically, when the data distribution is a mixture $p(x)=\pi p_{1}(x)+(1-\pi)p_{2}(x)$, where $p_{1}(x)$ and $p_{2}(x)$ are distinct and largely disjoint distributions, the gradient of the log probability density, $\nabla_{x}\log p(x)$, becomes problematic. In regions where $p_{1}(x)$ dominates, the score is driven solely by $p_{1}(x)$, and similarly for $p_{2}(x)$. Consequently, ULA, which relies on these score gradients, fails to correctly sample from the mixture distribution as it does not properly account for the mixing proportions $\pi$ and $1-\pi$. This leads to an incorrect estimation of the relative densities between the modes. The faster mixing time of MALA compared to Langevin has been well-established in the literature as in (Dwivedi et al., 2019; Wu et al., 2022).
To illustrate this, we conduct experiments to generate a mixture of Gaussian distributions with $p_{1}=\mathcal{N}\left((5,5)^{\top},I\right)$ and $p_{2}=\mathcal{N}\left((-5,-5)^{\top},I\right)$, with $\pi=0.8$. As can be observed in Figure 1, ULA produces samples that misrepresent the relative weights of the modes. On the other hand, we observe that the original MH algorithms significantly alleviate this problem, generating samples with mode weights close to the truth. This leads to our Score-Based MH framework, which facilitates the application of MH algorithms when only score functions are available.
@@
4 Algorithm
4.1 Acceptance Loss Function
Ideal Acceptance Loss. We begin with the ideal scenario where the true score function $\nabla\log p$ is known. In this scenario, we propose the following loss function
$\mathcal{L}_{i}(a)$ $=\mathbb{E}\bigg{[}\left\|\nabla\log a(X^{\prime},X)-\nabla\log a(X,X^{\prime})\right.$
@@ 4.1 Acceptance Loss Function
with $X\sim p$ and $X^{\prime}\sim q$, and the optimization objective is to learn an acceptance function $a^{*}\in\mathcal{F}$, where $\mathcal{F}$ is a set of hypotheses, such that
$a^{*}=\arg\min_{a\in\mathcal{F}}\mathcal{L}_{i}(a).$
@@
(a) Original samples.
(b) ULA.
(c) Score RW.
(d) Score MALA
(e) Score pCN.
Figure 5: Comparison of different methods on the Pinwheel dataset.
Practical Acceptance Loss. Typically, only samples from the data-generating process are observed, and the true $\nabla\log p$ is unknown. However, score parameterization through a neural network has proven successful in practice. We assume that we have an approximation of the score function previously estimated through a model $\tilde{s}\left(x\right)\approx\nabla\log p\left(x\right)$. Note that this is often the case in practice where an estimated score function $\tilde{s}$ is used in conjunction with ULA for sampling. In this scenario, the proposed practical loss function is
$\mathcal{L}_{p}(a)$ $=\mathbb{E}\bigg{[}\left\|\nabla\log a(X^{\prime},X)-\nabla\log a(X,X^{\prime})\right.$
$\quad-\left.\tilde{s}(X^{\prime})+\tilde{s}(X)-\nabla\log q(X\mid X^{\prime})\right.$
$\hskip 80.0pt\left.+\nabla\log q(X^{\prime}\mid X)\right\|^{2}\bigg{]},$ (5)
with $X\sim p$ and $X^{\prime}\sim q$, and the optimization objective is to learn an acceptance function $\tilde{a}\in\mathcal{F}$ where
$\tilde{a}=\arg\min_{a\in\mathcal{F}}\mathcal{L}_{p}(a).$
Illustrative Example.
We present the sampling results from a Gaussian distribution $\mathcal{N}(3,2)$ using the original Random Walk Metropolis-Hastings Algorithm, as shown in Figure 2(a). Next, we show the results of training with the ideal loss function for the Score-based RW in Figure 2(b), followed by the results using the practical loss function for Score-based RW in Figure 2(c).
4.2 Entropy Regularization

### v2

@@ Abstract
In this paper, we introduce a new approach for integrating score-based models with the Metropolis-Hastings algorithm. Score-based diffusion models have emerged as state-of-the-art methods for generative modeling, achieving remarkable performance in accurately learning the score function from data. However, these models lack an explicit energy function, making the Metropolis-Hastings adjustment step inaccessible. Consequently, the unadjusted Langevin algorithm is often used for sampling using estimated score functions. The lack of an energy function prevents the application of the Metropolis-adjusted Langevin algorithm and other Metropolis-Hastings methods, restricting the use of acceptance-function-based algorithms. We address this limitation by introducing a new loss function based on the detailed balance condition, allowing the estimation of the Metropolis-Hastings acceptance probabilities given a learned score function. We demonstrate the effectiveness of the proposed method for various scenarios, including sampling from score-based diffusion models and heavy-tail distributions.
@@ 1 Introduction
Sampling from probability distributions with access only to samples is a longstanding challenge that has received significant attention in machine learning and statistics (Goodfellow et al., 2020; Kingma, 2013; Ho et al., 2020; Dinh et al., 2016; Song et al., 2021). In particular, score-based models have emerged as a powerful class of generative algorithms, underpinning state-of-the-art (SOTA) models such as DALL·E 2 and Stable Diffusion (Rombach et al., 2022). Given samples from a target distribution, these models first estimate the score function —the gradient of the log-density of the data distribution (Hyvärinen, 2005)—using techniques like sliced score matching (Song et al., 2020) and denoising score matching (Vincent, 2011; Song and Ermon, 2019; Reu et al., ). Subsequently, they leverage the learned score function and score-based sampling methods, primarily the Unadjusted Langevin Algorithm (ULA), to sample from the target distribution (Grenander and Miller, 1994; Roberts and Tweedie, 1996).
@@
In parallel, a wealth of literature focuses on sampling from densities known only up to a normalizing constant (Hastings, 1970; Wang et al., 2024). A particular instance of this is the Metropolis-Hastings (MH) method, which provides a condition for accepting or rejecting new samples based on knowledge of a function proportional to the density (Robert et al., 2004). The MH method construct a Markov chain based on a proposal distribution and an acceptance function, ensuring convergence when the detailed balance condition is satisfied. Various MH algorithms have been introduced, defined by their choice of proposal distributions, such as the Random Walk Metropolis (RW) algorithm (Metropolis et al., 1953), the Metropolis-adjusted Langevin algorithm (MALA) (Roberts and Tweedie, 1996; Roberts and Stramer, 2002), and the preconditioned Crank-Nicolson (pCN) algorithm (Hairer et al., 2014). While these methods are theoretically well-grounded, they require knowledge of the unnormalized density, making them impractical when only samples or an estimated score function are available.
@@
A natural observation arises when considering score-based modeling: despite the extensive research on MH sampling, its integration with score-based methods remains largely unexplored. The key technical challenge lies in the absence of an unnormalized density when only the score function (or its estimation) is available. This work proposes a unifying framework that bridges this gap, enabling the application of MH algorithms using only samples and a learned score function. Consequently, this framework can leverage the strengths of both MH and score-based methods. Our analysis is motivated by two key questions:
@@
A). Can we estimate the acceptance function of an MH algorithm solely from an (estimated) score function and samples? B). Does including an MH adjustment step improve sample quality for score-based models?
@@
We answer these questions affirmatively by developing an approach to estimate an acceptance function from data. Our contributions are as follows:
  • We introduce Score Balance Matching, a method for estimating the acceptance function in the MH algorithm using only an estimated score function.
  • We use the learned acceptance function to incorporate MH steps into score-based samplers like ULA, enhancing the quality and diversity of generated samples.
  • We evaluate our method’s effectiveness and its robustness to hyperparameters by comparing it to ULA across diverse scenarios.
@@ 2 Related Work
A significant area of related work focuses on estimating a score function and developing efficient sampling strategies based on it. While most approaches aim to accelerate convergence to the target distribution (e.g. through improved numerical integrators), we consider our work as a parallel approach, incorporating an acceptance function to refine samples. This work lies at the intersection of the literature on score-based models and Metropolis-Hastings methods. We first review recent advances in both areas.
@@
Many successful generative models sample from an estimated score function (Song and Ermon, 2019; Song et al., 2021). These methods are trained using by minimizing the Fisher divergence between the true and predicted score (Hyvärinen, 2005). A significant research effort is concerned with understanding the efficacy of these sampling algorithms, with some papers investigating the asymptotic error of score-based diffusion models, as in (Chen et al., 2023; Zhang et al., 2024). Different sampling schemes have been proposed to improve sampling speed and quality (Dockhorn et al., 2021; Lu et al., 2022; Chen et al., 2024). These methods have relied heavily on various score-based samplers such as ULA or the reverse-time SDE of a diffusion process (Song et al., 2021).
Metropolis-Hastings.
Metropolis-Hastings algorithms are a class of MCMC algorithms that use an acceptance function to sample from target distributions (Metropolis et al., 1953; Hastings, 1970). These algorithms generate a sequence of samples by proposing a candidate state from a proposal distribution and accepting or rejecting the candidate based on an acceptance criterion that ensures convergence to the target distribution. The efficiency of MH algorithms heavily depends on the choice of the proposal distribution, and significant research effort has focused on designing efficient proposal mechanisms (Rosenthal et al., 2011; Song et al., 2017; Titsias and Dellaportas, 2019; Davies et al., 2023; Lew et al., 2023).
Recent developments in MH algorithms have explored the use of adaptive proposals to improve convergence rates and reduce autocorrelation in the generated samples (Andrieu and Thoms, 2008; Haario et al., 2001). These adaptive MH methods adjust the proposal distribution during the sampling process to better capture the geometry of the target distribution, leading to faster and more accurate sampling (Roberts and Rosenthal, 2009; Brooks et al., 2011; Hirt et al., 2021; Biron-Lattes et al., 2024).
(a) Original samples, weights $(0.8,0.2)$.
@@ 3 Score-Based Metropolis-Hastings
In this section, we first review score matching and MH algorithms in Sections 3.1 and 3.2 before we introduce Score Balance Matching in Section 3.3, a novel approach for learning the acceptance function from the score function. Let $X\sim p$ be a random variable, with support $\mathrm{supp}(X)=\mathcal{X}\subset\mathbb{R}^{d}$ and $p$ be its probability density function. We assume that $p$ is unknown, and only samples of $X$ are observed. Denote these observed samples by $\{x^{\left(i\right)}\}_{i=1}^{N}$, where $N$ is the number of observed samples.
3.1 Score Matching
Score-based models are grounded in score matching (Hyvärinen, 2005), which minimizes the Fisher divergence between the estimated and true gradients of the log data distribution, defined as follows.
Definition 1 (Fisher Divergence).
The Fisher divergence $D_{\nabla}(p_{1}\|p_{2})$ between two probability distributions $p_{1}$ and $p_{2}$ is defined as:
$D_{\nabla}(p_{1}\|p_{2})=\mathbb{E}_{x\sim p_{1}}\left[\|\nabla_{x}\log p_{1}(x)-\nabla_{x}\log p_{2}(x)\|^{2}\right].$
Therefore, given a class of hypothesis function $\mathcal{S}\subset\{s:\mathcal{X}\rightarrow\mathbb{R}^{d}\}$, score matching aims to find $s^{*}\in\mathcal{S}$ that minimizes the Fisher divergence to the data distribution $p$:
$s^{*}\in\operatorname*{arg\,min}_{s\in\mathcal{S}}\mathbb{E}_{x\sim p}\left[\|\nabla_{x}\log p(x)-s(x)\|^{2}\right]$ (1)
Since the score function is not directly observed, Hyvärinen proved that minimizing in the loss funcition in Equation 1 is equivalent to minimizing the following loss function:
$\mathbb{E}_{x\sim p}\left[\frac{1}{2}\|s(x)\|^{2}+\operatorname{tr}\left(\nabla_{x}s(x)\right)\right].$ (2)
Several variants of this loss have been proposed to circumvent the cumbersome estimation of the Hessian term in Equation 2, including sliced score matching and denoising score matching, defined as follows:
Definition 2 (Sliced Score Matching).
Sliced score matching (SSM) projects the score function onto random directions $v\sim\mathcal{N}(0,I_{d})$:
$\mathbb{E}_{x\sim p,v\sim\mathcal{N}(0,I_{d})}\left[\frac{1}{2}\|s(x)\|^{2}+v^{\top}\nabla_{x}s(x)v\right].$ (3)
Definition 3 (Denoising Score Matching).
Given a noise distribution $p_{\sigma}(\tilde{x}|x)$, the loss function is:
$\mathbb{E}_{x\sim p,\tilde{x}\sim p_{\sigma}(\tilde{x}|x)}\left[\|\nabla_{\tilde{x}}\log p_{\sigma}(\tilde{x}|x)-s(\tilde{x})\|^{2}\right].$ (4)
3.2 Metropolis-Hastings
MH involves an acceptance function $a:\mathcal{X}\times\mathcal{X}\rightarrow[0,1]$, which defines the probability $a(x^{\prime},x)$ of transitioning from the current $x\in\mathcal{X}$ to a proposal $x^{\prime}\in\mathcal{X}$. The common choice of the acceptance function is given by:
@@ 3.2 Metropolis-Hastings
Note that for computing $a(x^{\prime},x_{t})$ in Equation (5), a function proportional to $p(x)$ is needed.
@@
Although the acceptance function is often defined by Equation (5), a sufficient condition for ensuring convergence to the target distribution is that the acceptance ratio satisfies the detailed balance condition:
@@
In particular, any acceptance function satisfying the detailed balance condition above ensures that the Markov chain has the target distribution $p(x)$ as its unique stationary distribution. Note that the standard acceptance function in Equation (5) is one specific solution that satisfies this condition.
@@
3.3 Score Balance Matching
We now introduce our new objective function called Score Balance Matching (SBM). Let $q$ be a proposal distribution and let $\mathcal{A}$ be the set of valid acceptance functions for a given proposal $q$ and target $p$.
[…]
The SBM objective is defined as follows:
$\mathcal{L}(a)$ $=\mathbb{E}_{x\sim p,x^{\prime}\sim q(\cdot|x)}\bigg{[}\left\|\nabla\log a(x^{\prime},x)-\nabla\log a(x,x^{\prime})\right.$
@@ 3.3 Score Balance Matching
where $\nabla$ denotes the gradient with respect to $x,x^{\prime}\in\mathbb{R}^{d}$, i.e., $\nabla=\nabla_{x,x^{\prime}}$. The terms $\nabla\log p(x)$ and $\nabla\log p(x^{\prime})$ represent the score functions at $x$ and $x^{\prime}$, which can be approximated with a score model (e.g., obtained from score matching). We can express these score functions as:
and similarly,
Proposition 1.
We have that, for every function $a:\mathcal{X}\times\mathcal{X}\rightarrow[0,1]$
$\mathcal{L}(a)=0\iff a\in\mathcal{A}.$ (8)
Therefore, if we minimize SBM to zero, we are guaranteed to have a valid acceptance function. In practice, the acceptance function $a$ can be parameterized using an expressive class of functions and optimized to minimize SBM. Since we estimate the acceptance function from finite samples, we establish a generalization bound to quantify the approximation error when learning $a(x^{\prime},x)$ using a function class $\mathcal{F}$ (e.g., neural networks). We start by making the following assumptions.
Assumption 1.
$\mathcal{X}$ is compact.
Assumption 2.
$\nabla\log p$ and $\nabla\log q$ are continuous on $\mathcal{X}\times\mathcal{X}$.
Assumption 3.
$\forall a\in\mathcal{F},\nabla\log a$ is continuous on $\mathcal{X}\times\mathcal{X}$.
We note that Assumption 3 can be achieved by design. However, for many applications, Assumption 1 may not hold. To this end, we propose to employ the fact that the score function is scale invariant:
$\nabla_{x}\log p(x)=\nabla_{x}\log p^{*}(x),\quad\forall x\in\mathcal{X}.$
where, $\mathcal{X}\subset\mathbb{R}^{d}$ is chosen as a large compact subset of the support and $p^{*}$ is a truncated density:
$p^{*}(x)=\frac{p(x)}{Z}\mathbbm{1}\left(x\in\mathcal{X}\right),$
[…]
Proposition 2 (Finite-Sample Generalization Bound).
Let $\widehat{\mathcal{L}_{N}}(a)$ be the empirical SBM loss computed over $N$ i.i.d. training samples. Then, under Assumptions 12, and 3, we have that with probability at least $1-\delta$,
$\sup_{a\in\mathcal{F}}\left|\mathcal{L}(a)-\widehat{\mathcal{L}_{N}}(a)\right|\leq 2\mathcal{R}_{N}(\mathcal{F})+\mathcal{O}\left(\sqrt{\frac{\log(1/\delta)}{N}}\right),$ (9)
where $\mathcal{R}_{N}(\mathcal{F})$ is the empirical Rademacher complexity of class $\mathcal{F}$.
For the scope of this paper, we will instantiate our method on three representative MH algorithms, specified by their proposal functions: the Random Walk Metropolis-Hastings (RW), the Metropolis-Adjusted Langevin Algorithm (MALA), and the Preconditioned Crank-Nicolson (pCN) algorithm. We next briefly review these methods.
RW: The proposal function for RW is a Gaussian centered at the current state, i.e., $q(x^{\prime}|x)=\mathcal{N}\left(x,\sigma^{2}I\right)$.
MALA: The proposal distribution in MALA uses both the current state and the gradient of the log-probability, i.e., $q(x^{\prime}|x)=\mathcal{N}\left(x+\frac{\varepsilon^{2}}{2}\nabla\log p(x),\varepsilon^{2}I\right)$.
pCN: The proposal function for pCN is given by $x^{\prime}=\sqrt{1-\beta^{2}}x+\beta\xi$, where $\xi\sim\mathcal{N}(0,I)$.
When the acceptance function is learned using SBM, we refer to these methods as Score RW, Score MALA, and Score pCN, respectively. While we illustrate our results on these three MH algorithms, we note that the proposed framework is general and can be applied to other proposal functions. In Appendix A, we include another approach to approximate the acceptance function.
(a) Original samples.
(b) ULA.
(c) Score RW.
(d) Score MALA
(e) Score pCN.
Figure 3: Comparison of different methods on the Pinwheel dataset.
@@
3.4 Motivation
Our approach is motivated by three key challenges. First, traditional MH algorithms typically require the knowledge of the unnormalized density, while it is common in generative modeling to assume that only the score function can be accurately estimated. Second, ULA, widely used for sampling from score-based models, suffers from discretization errors that can lead to bias and slow-mixing. Consequently, by incorporating an MH adjustment step, we address a key limitation observed in ULA when sampling from distributions with multiple modes, particularly when these modes are separated by regions of low density. Specifically, when the data distribution is a mixture $p(x)=\pi p_{1}(x)+(1-\pi)p_{2}(x)$, where $p_{1}(x)$ and $p_{2}(x)$ are distinct and largely disjoint distributions, the gradient of the log probability density, $\nabla_{x}\log p(x)$, becomes misleading. In regions where $p_{1}(x)$ dominates, the score is driven solely by $p_{1}(x)$, and similarly for $p_{2}(x)$. Consequently, ULA, which relies on these score gradients, fails to correctly sample from the mixture distribution as it does not properly account for the mixing proportions $\pi$ and $1-\pi$. This leads to an incorrect estimation of the relative densities between the modes. The faster mixing time of MALA compared to Langevin has been well-established in the literature as in (Dwivedi et al., 2019; Wu et al., 2022). To illustrate this, we conduct experiments to generate a mixture of Gaussian distributions with $p_{1}=\mathcal{N}\left((5,5)^{\top},I\right)$ and $p_{2}=\mathcal{N}\left((-5,-5)^{\top},I\right)$, with $\pi=0.8$. Figure 1 shows that ULA produces samples that misrepresent the relative weights of the modes. In contrast, standard MH algorithms significantly alleviate this problem, generating samples with mode proportions that closely match the mixture weights.
4 Algorithm
In this section, we formulate the loss function and training algorithm for the acceptance function, which we model using neural networks.
Entropy Regularization.